What value of z* should be used to construct a 95% confidence interval of a population mean? Answer choices are rounded to the hundredths place.
Answer #1

At \(95 \%\) confidence level the \(z\) is,

\(\alpha=1-95 \%=1-0.95=0.05\)

\(\alpha / 2=0.05 / 2=0.025\)

\(\mathrm{Z}_{\alpha / 2}=\mathrm{Z}_{0.025}=1.960\)